OpenStudy (anonymous):

Solve 5^(x + 5) = 9^x

OpenStudy (anonymous):

x ≈ –3.63 x ≈ –0.28 x ≈ 0.07 x ≈ 13.69

OpenStudy (anonymous):

take the log of both sides: (x+9)log(5) = xlog(9) distribute: xlog5 + 9log5 = xlog9 bring the variable to 1 side: 9log5 = xlog9 - xlog5 combine like terms: 9log5 = x(log9 - log5) solve for x: x = (9log5)/(log9 - log5)

OpenStudy (anonymous):

and how do you solve for x?

OpenStudy (anonymous):

i just showed you in the last line... that is the exact value of x... if you want to see the value, put that expression into your calculator...

OpenStudy (anonymous):

oh okay thank you